OpenStudy (mw647):

Please help! Determine the changes in temperature of the Pb (ΔTPb). Data Table Pb mass of metal (g) 31.9812 volume of water (mL) 100 mass of water (g) 98.8 initial temperature of water (°C) 25.00 initial temperature of metal (°C) 200.00 max temp of water + metal (°C) 26.52

OpenStudy (anonymous):

Well the formula is q=m*s*T Q=joules or calories M=mass S=specific heat capacity (cal/g*celcius) T= change in temperature So you're looking for temperature which the formula is T=q/m*s

OpenStudy (anonymous):

What was the final temperature?